# Enhancing Neutrino Reconstruction in Water-Cherenkov Air Shower Arrays Using Multi-Photosensors

### v1

@@ Abstract
In this article, the potential of water Cherenkov detectors equipped with multi-PMT modules for precision neutrino direction reconstruction is demonstrated. By analyzing signal time traces with transformer-based models, significant improvements in angular resolution are achieved compared to previous designs with larger PMTs. These detectors enable the reconstruction of neutrino directions with resolutions of approximately $10^{\circ}$ in azimuth and $7^{\circ}$ in zenith for high-signal events. This design reduces saturation effects and enhances directional sensitivity, particularly for high-energy neutrinos. The results highlight the potential of WCD arrays as complementary tools for neutrino astronomy, particularly in the context of multimessenger observations of transient astrophysical sources. The nearly continuous operation and wide field of view of these detectors further enhance their suitability for real-time monitoring and alert generation.
@@ 1 Introduction
2 Detector concept and simulations
The WCD unit in this work is based on the Mercedes WCD design proposed in neutrinos_2024 ; wcd2022mercedes . The cylindrical detector retains its original dimensions (height $=1.7\,{\rm m}$, base area $\approx 12\,{\rm m^{2}}$) and features reflective interior walls to optimize Cherenkov light collection. Although the geometry remains unchanged, the photosensor configuration has been redesigned to improve the neutrino angular reconstruction performance.
@@ 2 Detector concept and simulations
Figure 2: Mean signal time trace for nearly vertical (black histogram: $\theta_{0}\geq 175^{\circ}$, $\phi_{0}\sim 180^{\circ}$) and inclined (red histogram: $\theta_{0}\leq 150^{\circ}$, $\phi_{0}\sim 180^{\circ}$) upward-going muons with an energy of $2\,\rm GeV$. The events are centered within a radius of $r\leq 0.5\,\rm m$. The PMTs were arranged in a hexagonal layout to represent the real orientation of each PMT in the multi-PMT module.
@@
For this study, a total of roughly 150 000150000150\,000150 000 upward-going $1\,\rm TeV$ muon-neutrinos were simulated. Only events exceeding a signal threshold of $100$ photoelectrons (p.e.) were processed, a threshold determined based on the typical signal produced by vertically crossing muons in the WCD. This cut ensures that the analysis is focused on events in which particles fully cross the detector, while effectively suppressing low-energy background sources neutrinos_2024 . After applying this cut, the dataset was divided into 44 1074410744\,10744 107 events for training, 11 0271102711\,02711 027 for validation, and 69 6526965269\,65269 652 for testing. The training set is used to optimize the model, the validation set for hyperparameter tuning and overfitting prevention, and the test set for unbiased performance evaluation.
@@
The analysis in this work is based on the analysis of signal time traces, which are constructed through a detailed simulation of the PMT electronics chain. For each photoelectron, a pulse is generated using reference time and amplitude distributions, and the resulting analog waveform is formed by summing these individual pulses [^1] The reference distributions for the PMT electronics simulation chain were taken from Electronics .. The waveform is then digitized by sampling its amplitude at regular intervals, determined by a sampling rate of 250 mega samples per second (MSPS). To account for potential misalignment with the master clock, the start of digitization is randomized within the time resolution. Before digitization, an offset of $0.8$ Least Significant Bit (LSB) is added to the signal. The digitization process employs a 12-bit Analog-to-Digital Converter (ADC) with a range of $[0,4095]$. In this step, the signal is truncated to the nearest integer and an offset of $400\,\rm LSB$ is added. The simulation includes PMT saturation effects, with pulses exceeding $\sim\!250\,\mathrm{p.e.}$ being clipped by the 12-bit ADC’s dynamic range.
To minimize the influence of electronics on the traces, a preprocessing step is applied after the PMT electronics simulation. First, a baseline of $402\,\rm LSB$ is subtracted from the signals to account for the offset and Gaussian noise introduced by the electronics. Second, the $T_{0}$ of the event is determined by identifying the first time bin in which the cumulative signal across all PMTs exceeds a threshold of $5\,\rm LSB$ above the baseline. This $T_{0}$ is then used to shift the time traces of all PMTs equally, aligning the event start to the beginning of the trace. In this way, the relative timing between signals in different PMTs is preserved, ensuring that temporal features such as the time differences between peaks across PMTs remain intact. Once $T_{0}$ is set, a time window of 32 ns after $T_{0}$ is used for further analysis, which is sufficient to capture the direct Cherenkov light and the first peak of reflected Cherenkov light wcd2022mercedes (see Figure 2).
@@ 3 Methodology for angular reconstruction
The transformer models were implemented using PyTorch, with hyperparameters such as the number of layers, attention heads, and feedforward dimensions optimized using the Optuna framework optuna2019 . The optimal configuration included 2 transformer layers, 8 attention heads, a model dimension of 32, and a feedforward dimension of 128. Furthermore, the model featured two dense layers with 64 and 16 neurons, respectively, using ReLU activation and dropout rates of 0.1. The Adam optimizer kingma2014adam was used with a learning rate of $0.001$, betas of $(0.9,0.999)$, and a learning rate decay of $10^{-8}$. Training was conducted over 1 00010001\,0001 000 epochs with a batch size of 1 02410241\,0241 024 and an early stopping tolerance of 50 epochs to avoid overfitting. To ensure reproducibility and efficient monitoring of the training process, the results of the experiments were systematically tracked and managed using MLflow MLflow2018 .
For the classification of the entry zone, a standard cross-entropy loss was used. For the angular reconstruction task, on the other hand, the loss function was specifically designed to handle the periodic nature of angles, as proposed in neutrinos_2024 . A circular regression loss function was used, defined as: ${\rm Loss}=\frac{1}{2N}\sum^{N}_{i=1}\left[1-\cos(y_{i}-\hat{y}_{i})\right]\in[0,1]$. Here, $N$ is the number of samples, $y_{i}$ is the true angle for the $i$-th sample, and $\hat{y}_{i}$ is the predicted angle. This formulation ensures accurate reconstruction within the range of 0 to 360 degrees by explicitly accounting for the periodicity of angles. The loss is normalized by $1/2N$ to scale it between 0 and 1.
@@ 4 Results
Figure 4: Histogram of the angular distance in $\phi$ between the true and predicted entry zones for neutrino events. An angular distance of zero denotes a correct classification of the entry zone.
@@
Figure 5: Comparison of the angular reconstruction efficiency as a function of the signal for the M1T1 (gray, neutrinos_2024 ), M3T1 (blue, neutrinos_2024 ), M1mT1s (dashed black line, this work), and M1mT1m (orange, this work) detectors. The top panel shows the angular reconstruction efficiency, while the bottom panel displays the number of events within each signal bin, normalized to the total number of events in the test dataset. Error bars represent the $68\%$ containment for each signal bin. Note that the M1T1 detector is not able to reconstruct $\phi$ due to the azimuthal symmetry of its PMTs.
@@
[…]
@@
To put these results in context, dedicated neutrino experiments such as IceCube and KM3NeT/ARCA achieve angular resolutions better than $1^{\circ}$ for track-like muon neutrino events in the TeV energy range. For cascade-like events, the resolutions are $>10^{\circ}$ for IceCube and $2^{\circ}-3^{\circ}$ for KM3NeT/ARCA km3net2025PeV ; KM3NeT2019sensitivity ; icecube2013EffMass ; icecube2023observation . Although the resolution of the M1mT1m detector is less precise than dedicated experiments, it is important to note that this WCD was primarily designed for gamma-ray detection. Despite this, the achieved resolution is competitive, making the M1mT1m detector a valuable complementary tool for neutrino astronomy, particularly for high-energy events where the neutrino flux is low and every detection is critical Ackermann:2022rqc ; Kurahashi:2022utm . It should also be noted that transient sources, such as gamma-ray bursts, active galactic nuclei, and neutron star mergers, are expected to produce neutrinos in bursts waxman1997high ; ioka2005tev ; abbasi2011neutrino . In such cases, the temporal clustering of neutrino events, combined with their spatial coincidence, can provide strong evidence of their astrophysical origin, even if individual angular resolutions are limited icecube2013EffMass .
@@ 5 Conclusions
In this work, we demonstrated the potential of WCDs equipped with multi-PMT modules for neutrino direction reconstruction. The reconstruction is based on the analysis of signal time traces using transformer-based models. The multi-PMT-based WCD, featuring seven 3-inch PMTs arranged in a hexagonal pattern, provides enhanced directional sensitivity. This design enables the reconstruction of neutrino directions with resolutions of approximately $10^{\circ}$ in azimuth and $7^{\circ}$ in zenith for high-signal events, representing a substantial improvement over results obtained with WCDs using multiple larger PMTs.
@@
[…]

### v2

@@ Abstract
In this article, the potential of water Cherenkov detectors equipped with multi-PMT modules for background-free upward-going neutrino detection and improved direction reconstruction is demonstrated. By analyzing signal time traces with transformer-based models, significant improvements in angular resolution are achieved compared to previous designs with larger PMTs. These detectors enable the reconstruction of neutrino directions with resolutions of approximately $10^{\circ}$ in azimuth and $7^{\circ}$ in zenith for high-signal events, corresponding to an overall opening angle of approximately $10^{\circ}$. This design reduces saturation effects and enhances directional sensitivity, particularly for high-energy neutrinos. The results highlight the potential of WCD arrays as complementary tools for neutrino astronomy, particularly in the context of multimessenger observations of transient astrophysical sources. The nearly continuous operation and wide field of view of these detectors further enhance their suitability for real-time monitoring and alert generation.
@@ 1 Introduction
2 Detector concept and simulations for angular reconstruction
The WCD unit in this work is based on the Mercedes WCD design proposed in neutrinos_2024 ; wcd2022mercedes . The cylindrical detector retains its original dimensions (height $1.7\,{\rm m}$, base area $\approx 12\,{\rm m^{2}}$) and features reflective interior walls to optimize Cherenkov light collection. Although the geometry remains unchanged, the photosensor configuration has been redesigned to improve the neutrino angular reconstruction performance.
@@ 2 Detector concept and simulations for angular reconstruction
Figure 2: Preprocessed mean signal time trace for nearly vertical (black histogram: $\theta_{0}\geq 175^{\circ}$, $\phi_{0}\sim 180^{\circ}$) and inclined (red histogram: $\theta_{0}\leq 150^{\circ}$, $\phi_{0}\sim 180^{\circ}$) upward-going muons with an energy of $2\,\rm GeV$. The events are centered within a radius of $r\leq 0.5\,\rm m$. The waveform resulting from the PMT electronics simulation, originally in LSB units, has been preprocessed to determine $T_{0}$, the baseline of $402\,\rm LSB$ has been subtracted, and it has been normalised by the gain $G=15\,\rm LSB/p.e.$ The PMTs were arranged in a hexagonal layout to represent the real orientation of each PMT in the multi-PMT module.
@@
For this study, a total of roughly 150 000150\,000 upward-going $1\,\rm TeV$ muon-neutrinos were simulated. Only events exceeding a signal threshold of $100$ photoelectrons (p.e.) were processed, a threshold determined based on the typical signal produced by vertically crossing muons in the WCD. This cut ensures that the analysis is focused on events in which particles fully cross the detector, while effectively suppressing low-energy background sources neutrinos_2024 . After applying this cut, the dataset was divided into 44 10744\,107 events for training, 11 02711\,027 for validation, and 69 65269\,652 for testing. The training set is used to optimize the model, the validation set for hyperparameter tuning and overfitting prevention, and the test set for unbiased performance evaluation.
@@
The analysis in this work is based on the analysis of signal time traces, which are constructed through a detailed simulation of the PMT electronics chain. For each photoelectron, a pulse is generated using reference time and amplitude distributions, and the resulting analog waveform is formed by summing these individual pulses [^1] The reference distributions for the PMT electronics simulation chain were taken from Electronics .. In addition, signal noise is included based on measurements from the experimental setup described in  Electronics . This noise was found to follow a Gaussian distribution with a standard deviation of approximately $\sigma\sim 0.8\,\mathrm{LSB}$, which is added to the analog signal before digitization.
The waveform is then digitized by sampling its amplitude at regular intervals, determined by a sampling rate of 250 mega samples per second (MSPS). To account for the arbitrary phase offset between the signal generation and the ADC sampling clock, the start of the digitization process is randomized within one sampling interval. This simulates the lack of synchronization with the master clock and introduces a realistic phase jitter in the digitized waveform. Specifically, a random time shift uniformly distributed in the interval $0,\,1/f_{\mathrm{sample}}]$ is applied, where $f_{\mathrm{sample}}$ is the sampling frequency, which is $250\,\rm MSPS$. This approach reproduces the variations expected from unsynchronized sampling and enhances the realism of the simulation by introducing up to 4 ns uncertainty in the alignment between the digitized signal and its true onset. The digitization process employs a 12-bit Analog-to-Digital Converter (ADC) with a range of $[0,4095]$. In this step, the signal is truncated to the nearest integer and an offset of $400\,\rm LSB$ is added. The simulation includes PMT saturation effects, with pulses exceeding $\sim\!250\,\mathrm{p.e.}$ being clipped by the 12-bit ADC’s dynamic range.
To minimize the influence of electronics on the traces, a preprocessing step is applied after the PMT electronics simulation. First, a baseline of $402\,\rm LSB$ is subtracted from the signals to account for the offset and Gaussian noise introduced by the electronics. Second, the $T_{0}$ of the event is determined by identifying the first time bin in which the cumulative signal across all PMTs exceeds a threshold of $5\,\rm LSB$ above the baseline. This $T_{0}$ is then used to shift the time traces of all PMTs equally, aligning the event start to the beginning of the trace. In this way, the relative timing between signals in different PMTs is preserved, ensuring that temporal features, such as the time differences between peaks across PMTs, remain intact. Once $T_{0}$ is set, the signal is divided by the electronics gain, $G=15\,\rm LSB/p.e.$ , and a time window of 32 ns after $T_{0}$ is used for further analysis, which is sufficient to capture the direct Cherenkov light and the first peak of reflected Cherenkov light wcd2022mercedes (see Figure 2).
@@
3 Neutrino effective mass
To demonstrate the capability of the M1mT1m detector to discriminate particles induced by upward-going neutrinos from those produced by cosmic rays, the analysis pipeline developed in neutrinos_2024 was applied to this new detector configuration.
[…]
The second dataset, used for the calculation of the neutrino effective mass, contains approximately $2.5\cdot 10^{5}$ upward-going muon and electron neutrino events per simulation bin. These events are generated with fixed energies of $\{10\,\rm GeV,1\,\rm TeV,100\,\rm TeV\}$ and various zenith angles: $\{0^{\circ},15^{\circ},30^{\circ},40^{\circ}\}$.
The identification chain begins with basic cuts on the signal time traces, designed to suppress background contamination. For simplicity, the summed response of the small PMTs in each multi-PMT module was used to approximate the performance of the single 8-inch PMT employed in the original M1T1 configuration. Although a dedicated set of cuts adapted to the geometry of the multi-PMT module could enhance performance, this lies beyond the scope of this work.
The applied selection criteria require: first, the maximum signal in the upper multi-PMT to occur in the first time bin, and second, the maximum in the lower multi-PMT to fall between the second and fourth bins. These conditions correspond to the direct detection of Cherenkov light followed by its reflection within the WCD, with a minimum time delay of $8\,\rm ns$, based on the $4\,\rm ns$ sampling rate and the $1.7\,\rm m$ vertical height of the water volume. The chosen maximum positions are visible in the mean signal time traces of nearly vertical events in Figure 2. These signal cuts reject approximately $85\%$ of the background while preserving about $50\%$ of the signal events. These results are obtained under a conservative assumption for the detector electronics, where the PMT sampling rate is $250\,\rm MSPS$ (corresponding to $4\,\rm ns$ time bins), matching the minimum time resolution required to apply the signal cuts effectively. Improved PMT electronics with higher sampling rates (i.e., thinner time bins) could potentially enhance the background rejection capability further by allowing a more precise characterization of the signal timing structure neutrinos_2024 .
In the final step of the analysis, a CNN is used to classify the surviving events. The CNN was trained to produce outputs of 1-1 for upward-going events (originating from negative $z$) and +1+1 for downward-going events (positive $z$). As shown in Figure 3, all background particles yield CNN outputs above 0.5, confirming that none survive both the signal and CNN cuts. In contrast, a fraction of the neutrino-induced events are correctly identified, yielding outputs near 1-1.
Figure 3: Histogram of the CNN output for neutrino identification using the M1mT1m detector. Only the events passing the signal cuts were used to produce this plot.
To quantify the overall efficiency of the method, we estimate the neutrino effective mass for a WCD array composed of 5 4615\,461 M1mT1m detectors. The effective mass depends on the neutrino’s direction $(\theta,\phi)$, energy ($E_{\nu}$), the number of WCDs in the EAS array (N), and interaction point $(x,y,D)$, where $D$ is the slant depth (in $\rm g\,cm^{-2}$}) from the WCD station to the neutrino. Assuming a uniform distribution in $\phi$, the effective mass can be computed either for an isotropic flux using Equation (1).
$M_{\text{eff}}\ (E_{\nu})$ $=\int\dfrac{dM_{\text{eff}}}{d\theta}(\theta,E_{\nu})\ {\rm d}\theta$ (1)
$=2\pi N\int\sin{\theta}\ \varepsilon(x,y,D,\theta,E_{\nu})\ {\rm d}x\ {\rm d}y\ {\rm d}D\ {\rm d}\theta\ {\rm[g]}$
In Figure 4, the resulting effective mass is presented and compared with the one of the M1T1 detector neutrinos_2024 . This result demonstrates that the M1mT1m detector can provide effective background-free neutrino identification. For electron neutrinos, the effective mass is on the order of $10^{4}\,\rm tons$, comparable to that of the Super-Kamiokande experiment. In the case of muon neutrinos, the effective mass increases more significantly with energy, as muons can traverse longer distances through the ground before reaching the detector. Specifically, it reaches approximately $10^{4}\,\rm tons$ at $10\,\rm GeV$—again comparable to Super-Kamiokande—scales up to $10^{5}\,\rm tons$ at $1\,\rm TeV$, similar to the water mass of Hyper-Kamiokande, and ultimately approaches $10^{7}\,\rm tons$ at $100\,\rm TeV$, closer to the performance of IceCube for High-Energy Starting Events.
The performance degradation compared to the M1T1 detector presented in neutrinos_2024 is mainly attributed to the lower sampling rate used in the electronics simulation, rather than to the detector concept itself: $250\,\rm MSPS$ for the M1mT1m configuration versus $1\,\rm GSPS$ in the previous study. In particular, this reduced sampling rate has a significant impact on the efficiency of the signal cuts. While the $1\,\rm GSPS$ setup achieved a background rejection rate of approximately $99\%$, similar signal cuts only reach about $85\%$ efficiency with the $250\,\rm MSPS$ configuration adopted here.
This result underscores the critical role of the front-end electronics in enabling effective background rejection. In the present study, a conservative, worst-case scenario is considered, where the PMT sampling rate is equal to the time resolution required to apply the signal cuts described earlier. Improving the sampling rate could thus enhance the detector’s performance in terms of background suppression.
Figure 4: Neutrino effective mass obtained for a diffuse flux of neutrinos using the M1mT1m detector with a $250\,\rm MSPS$ sampling rate ($4\,\rm ns$ time bins), compared with the results from our previous study using the M1T1 detector with a $1\,\rm GSPS$ sampling rate neutrinos_2024 ($1\,\rm ns$ time bins) both using a EAS array with 5 4615\,461 WCDs. The results are compared with the effective masses of the following experiments: Super-Kamiokande SuperKamiokande , Hyper-Kamiokande HyperKamiokande , DeepCore online filter for $\nu_{\mu}$ (taken from Fig. 2 of reference DeepCore2016 ), KM3NeT/ORCA (taken from Fig. 1 of reference KM3NeT2022EffMass ), IceCube (derived from Figures 2 and 5 of reference icecube2023observation ), and IceCube $\nu_{e}$ and $\nu_{\mu}$ High-Energy Starting Events (HESE) (taken from Fig. 7B of reference icecube2013EffMass ). A shaded area between the fiducial and total water volume of the Kamiokande detectors was added.
@@ 4 Methodology for angular reconstruction
The transformer models were implemented using PyTorch, with hyperparameters such as the number of layers, attention heads, and feedforward dimensions optimized using the Optuna framework optuna2019 . The optimal configuration included 2 transformer layers, 8 attention heads, a model dimension of 32, and a feedforward dimension of 128. Furthermore, the model featured two dense layers with 64 and 16 neurons, respectively, using ReLU activation and dropout rates of 0.1. The Adam optimizer kingma2014adam was used with a learning rate of $0.001$, betas of $(0.9,0.999)$, and a learning rate decay of $10^{-8}$. Training was conducted over 1 0001\,000 epochs with a batch size of 1 0241\,024 and an early stopping tolerance of 50 epochs to avoid overfitting. To ensure reproducibility and efficient monitoring of the training process, the results of the experiments were systematically tracked and managed using MLflow MLflow2018 .
For the classification of the entry zone, a standard cross-entropy loss was used. For the angular reconstruction task, on the other hand, the loss function was specifically designed to handle the periodic nature of angles, as proposed in neutrinos_2024 . A circular regression loss function was used, defined in Equation 2, where $N$ is the number of samples, $y_{i}$ is the true angle for the $i$-th sample, and $\hat{y}_{i}$ is the predicted angle. This formulation ensures accurate reconstruction within the range of 0 to 360 degrees by explicitly accounting for the periodicity of angles. The loss is normalized by $1/2N$ to scale it between 0 and 1.
${\rm Loss}=\frac{1}{2N}\sum^{N}_{i=1}\left[1-\cos(y_{i}-\hat{y}_{i})\right]\in[0,1]$ (2)
@@ 5 Results
(a) Histogram of the angular distance in $\phi$ between the true and predicted entry zones for neutrino events. An angular distance of zero denotes a correct classification of the entry zone.
(b) Visualization of the entry points on the WCD surface, colored according to the reconstructed entry zone predicted by the transformer. The background colors represent the entry zones, which follow an anticlockwise order starting with Zone 0 centered at $0^{\circ}$ (red color).
Figure 6: Performance of the entry zone reconstruction.
@@
Figure 7: Comparison of the angular reconstruction efficiency with respect to the true neutrino direction as a function of the signal for the M1T1 (gray, neutrinos_2024 ), M3T1 (blue, neutrinos_2024 ), M1mT1s (dashed black line, this work), and M1mT1m (orange, this work) detectors. The top panel shows the angular reconstruction efficiency, while the bottom panel displays the number of events within each signal bin, normalized to the total number of events in the test dataset. Error bars represent the $68\%$ containment for each signal bin. Note that the M1T1 detector is not able to reconstruct $\phi$ due to the azimuthal symmetry of its PMTs.
@@
The multi-PMT-based WCDs demonstrate a significant improvement over those using single larger PMTs, in particular for the reconstruction of the azimuthal angle. While the M1T1 is not sensitive to the azimuthal direction of the neutrino, the M1mT1m design improves the M3T1 detector angular resolution by a factor of two, being $10^{\circ}$ for high-signal events. This enhancement is largely due to the distribution of the PMTs: in the M3T1 detector, the PMTs were positioned close to the WCD lateral wall in a $120^{\circ}$ star configuration, while in the M1mT1m detector, the peripheral PMTs are separated by $60^{\circ}$. Moreover, the use of multiple smaller PMTs may reduce saturation effects, improving the reconstruction efficiency at the highest energies. It is important to highlight that, for events with $S\geq 1\,000\,\rm p.e.$, the M1mT1s exhibits a similar performance as the M1mT1m. This result indicates that the upper multi-PMT module is essential to reconstruct low-signal events, effectively using direct Cherenkov light, while the patterns found on the reflected light captured on the bottom multi-PMT are enough to achieve an accurate angular reconstruction in high-signal events.
@@
To put these results in context, the opening angle between the reconstructed and true neutrino directions in this analysis is approximately $10^{\circ}$. In comparison, dedicated neutrino experiments such as IceCube and KM3NeT/ARCA achieve angular resolutions better than $1^{\circ}$ for track-like muon neutrino events in the TeV energy range. For cascade-like events, the resolutions are typically greater than $10^{\circ}$ for IceCube and around $2^{\circ}$$3^{\circ}$ for KM3NeT/ARCA km3net2025PeV ; KM3NeT2019sensitivity ; icecube2013EffMass ; icecube2023observation . Although the resolution of the M1mT1m detector is less precise than dedicated experiments, it is important to note that this WCD was primarily designed for gamma-ray detection. Despite this, the achieved resolution is competitive, making the M1mT1m detector a valuable complementary tool for neutrino astronomy, particularly for high-energy events where the neutrino flux is low and every detection is critical Ackermann:2022rqc ; Kurahashi:2022utm . It should also be noted that transient sources, such as gamma-ray bursts, active galactic nuclei, and neutron star mergers, are expected to produce neutrinos in bursts waxman1997high ; ioka2005tev ; abbasi2011neutrino . In such cases, the temporal clustering of neutrino events, combined with their spatial coincidence, can provide strong evidence of their astrophysical origin, even if individual angular resolutions are limited icecube2013EffMass .
@@ 6 Conclusions
In this work, we demonstrated the potential of WCDs equipped with multi-PMT modules for background-free neutrino detection and direction reconstruction, provided that the sampling rate of the PMT electronics is $250\,\rm MSPS$ or better. The reconstruction is based on the analysis of signal time traces using transformer-based models. The multi-PMT-based WCD, featuring seven 3-inch PMTs arranged in a hexagonal pattern, provides enhanced directional sensitivity. This design enables the reconstruction of neutrino directions with angular resolutions of approximately $10^{\circ}$ in azimuth and $7^{\circ}$ in zenith for high-signal events, corresponding to a mean opening angle of about $10^{\circ}$ with respect to the true neutrino direction. This result represents a substantial improvement over results obtained with WCDs using multiple larger PMTs.
@@
A cost-effective alternative configuration was also evaluated, featuring a multi-PMT module at the bottom and a single 3-inch PMT at the top, demonstrating performance comparable to the dual multi-PMT design for high-signal events ($S\geq 1\,000\,\rm p.e.$).